# Xinwei Ma Jingshen Wang Waverly Wei

### v2

@@ 1.1 Motivation and contribution
In clinical trials, social science field experiments, and A/B tests, participants often enroll sequentially, and responses to treatments can vary due to individual-specific characteristics. In these contexts, covariate-adjusted response-adaptive (CARA) designs have gained increasing popularity: by dynamically adjusting treatment assignments based on accumulated covariates and outcome information during the experiment, CARA designs can be adopted to optimize treatment allocation to improve statistical power (Hahn, Hirano and Karlan, 2011; Hu, Zhu and Hu, 2015; Blackwell, Pashley and Valentino, 2023), reduce participant exposure to less effective treatments and enhance overall welfare (Hu and Rosenberger, 2006; Hu, Zhu and Hu, 2015; Wei, Ma and Wang, 2024, 2025), and maintain statistical validity for analyzing experimental data (Zhang, Hu, Cheung and Chan, 2007; Hu and Hu, 2012; Bugni, Canay and Shaikh, 2018; Offer-Westort, Coppock and Green, 2021; Bai, Romano and Shaikh, 2022; Robertson, Lee, López-Kolkovska and Villar, 2023; Zhao, 2023; Wei, Ma and Wang, 2025; Bibaut and Kallus, 2025).
@@
As an example, we revisit the study by Fahey, Njau, Katabaro, Mfaume, Ulenga, Mwenda et al. (2020), a randomized experiment conducted in the Shinyanga region of Tanzania between April 24 and December 14, 2018. The study aimed to assess the impact of cash incentives on retention in care and viral suppression among people living with HIV (PLHIV). A total of 530 PLHIV aged 18 or older were sequentially enrolled and randomly assigned to either a treatment group receiving cash incentives or a control group without any incentive. Viral load was measured six months after treatment assignment through blood draws. However, delays occurred, as the lab test required a separate clinical visit and some participants missed their scheduled appointments. We calculated the number of days delayed as the difference between the date of the viral load test and the six-month follow-up date, separately for two covariate strata defined by biological sex. From Figure 1, it is clear that the distribution of outcome delays varies significantly depending on treatment assignment and gender. Specifically, the treatment arm exhibits less delay, and within the treatment arm, the delayed response issue seems less pronounced among female participants. Overall, the empirical observation highlights the importance of accounting for delayed outcomes that are both arm- and covariate-specific.
@@
To our knowledge, although there has been prior work on developing valid statistical inference procedures in the presence of delayed responses in CARA (Bai, Hu and Rosenberger, 2002; Hu, Zhang, Cheung and Chan, 2008; Robertson, Lee, López-Kolkovska and Villar, 2023), no CARA designs has incorporated the delay mechanism when optimizing experimental objectives. In this paper, we address the challenges that response delays pose to covariate-adjusted response-adaptive (CARA) designs. We begin by examining whether existing optimal allocation strategies remain valid under delayed outcomes and, when they do not, derive the optimal allocation in such settings. We then develop methods to adapt CARA designs accordingly. Finally, we propose an efficient estimator that accounts for delayed responses at the end of the experiment. We elaborate on these contributions in more detail below.
@@
First, we demonstrate that the classical optimal allocation strategy used in existing CARA designs no longer provides optimal treatment assignment when accounting for arm- and covariate-dependent delays. Specifically, we first derive a semiparametric efficiency bound for treatment effect estimation in the presence of outcome delay (Theorem 1), which differ substantially from existing results in the literature (Hahn, 1998; Hirano, Imbens and Ridder, 2003; Cattaneo, 2010; Zhu and Zhu, 2023; Wu, Zheng, Zhang, Zhang and Wang, 2025). See Van der Vaart (2000) and Tsiatis (2006) for general discussions on semiparametric efficiency. This result allows us to show that the optimal delay mechanism involves a unique trajectory of treatment assignment probabilities that vary across different stages, resulting in treatment allocations with lower variance for power maximization, and enhanced failure reduction in welfare improvement designs. To our knowledge, this is the first design to explicitly explore delay mechanisms to optimize trial objectives, while existing literature focuses on conducting sequential tests (Hampson and Jennison, 2013), designing Bayesian adaptive designs where delay is independent of the arm and covariates (Lin, Thall and Yuan, 2020), or conducting statistical inference after experiments have concluded (Bai, Hu and Rosenberger, 2002; Hu, Zhang, Cheung and Chan, 2008).
@@
Second, since the delay mechanism is unknown a priori and remains only partially estimable during the experiment, we propose a fully forward-looking CARA design that sequentially updates treatment assignment probabilities to target a trial objective at the conclusion of the experiment. In this process, our algorithm offers several extrapolation strategies when estimating delay mechanism, allowing better alignment with the designer’s prior knowledge. We further discuss in detail two design objectives frequently adopted in current CARA literature: maximizing experiment power (Tymofyeyev, Rosenberger and Hu, 2007; Zhao, 2023) and improving overall participant welfare (or failure reduction) while maintaining statistical power constraints (Rosenberger, Stallard, Ivanova, Harper and Ricks, 2001; Robertson, Lee, López-Kolkovska and Villar, 2023). We justify the benefits of our proposed forward-looking CARA design through theoretical investigation (Section 4) and a synthetic case study (Section 5). Under the design objective of maximizing experiment power, our proposed forward-looking CARA design attains higher estimation efficiency compared to the Neyman allocation and complete randomization designs in the presence of delayed responses. Under the design objective of improving overall participant welfare, we demonstrate that our proposed forward-looking CARA design achieves additional failure reduction, leading to greater welfare improvement compared to both the ethical design and the complete randomization design while maintaining statistical power. For both design objectives, our proposed design can provide valid statistical inference, which is justified in Theorem 3.
@@ 2 Optimal delay-adjusted treatment allocation in CARA
From the above result, it is evident that when the delay mechanism of the response depends on the covariates and treatment status, the optimal treatment allocation rules in classical CARA designs must be modified to account for outcome delays. In what follows, we examine how the optimal treatment allocation rule is affected by the arm- and/or covariate-dependent delayed responses in two common experimental objectives: one aimed at enhancing statistical power (Tymofyeyev, Rosenberger and Hu, 2007), and the other focused on minimizing the expected number of failures while maintaining a minimum power requirement for binary outcomes (Rosenberger, Stallard, Ivanova, Harper and Ricks, 2001). In both design goals, the variance of the estimated treatment effects plays an essential role. To simplify the presentation, we focus on two-arm CARA designs with $a\in\{0,1\}$ in this paper. For multi-arm experiments, the design objective can be generalized using non-centrality parameters, as discussed in Lachin (1977) and Tymofyeyev, Rosenberger and Hu (2007).
@@
where $\delta$ ensures that the assignment probabilities are bounded away from 0 and 1, maintaining sufficient randomness in treatment assignments (Ma and Wang, 2020; Heiler and Kazak, 2021; Sasaki and Ura, 2022; Ma, Sasaki and Wang, 2025; Dorn, 2025). The objective function $\mathsf{V}(e_{1}(\cdot),\dots,e_{T}(\cdot))$ is a component of the semiparametric efficiency bound related to the propensity scores, and is defined as:
@@
Design objective 2: Optimal treatment allocation for failure reduction subject to a power constraint. This design aims to minimize the expected number of failures while satisfying a minimum power requirement for binary responses (Rosenberger, Stallard, Ivanova, Harper and Ricks, 2001). In this context, a treatment is considered to fail when the potential outcome equals one. The optimal allocation thus seeks to assign treatments based on the solution to the following optimization problem:
@@
and the optimal failure reduction rule of Rosenberger, Stallard, Ivanova, Harper and Ricks (2001):
@@ 3 Fully forward-looking CARA with delayed outcomes
The results provided in the previous section assume prior knowledge of the delay mechanism, and therefore, the discussed optimal treatment allocation corresponds to an infeasible oracle setting. In practice, the delay mechanism is rarely known and must be estimated as data are sequentially accumulated to optimize the experimental objective. In this section, we propose two delay-informed practical CARA designs tailored to two previously discussed experimental goals: improving statistical power (Tymofyeyev, Rosenberger and Hu, 2007) and minimizing the expected number of failures while meeting a minimum power requirement for binary responses (Rosenberger, Stallard, Ivanova, Harper and Ricks, 2001). We also provide tools for valid statistical inference after the experiment is finished.
@@
$\hat{\mathsf{V}}_{t}(e_{t+1}(\cdot),\dots,e_{T}(\cdot))$ $=\sum_{x\in\mathcal{X}}\hat{p}_{t}(x)\Bigg{[}\frac{\hat{\sigma}_{t}^{2}(x,1)}{\sum\limits_{\ell=1}^{t}\hat{r}_{\ell}\hat{\rho}_{t}(T-\ell|x,1)\hat{e}^{*}_{\ell}(1|x)+\sum\limits_{\ell=t+1}^{T}{r}_{\ell}\hat{\rho}_{t}(T-\ell|x,1)e_{\ell}(x)}$
$\qquad\qquad+\frac{\hat{\sigma}_{t}^{2}(x,0)}{\sum\limits_{\ell=1}^{t}\hat{r}_{\ell}\hat{\rho}_{t}(T-\ell|x,0)\hat{e}^{*}_{\ell}(0|x)+\sum\limits_{\ell=t+1}^{T}{r}_{\ell}\hat{\rho}_{t}(T-\ell|x,0)(1-e_{\ell}(x))}\Bigg{]}.$
@@
$\quad+\sum_{\ell=t+1}^{T}r_{l}\sum_{x\in\mathcal{X}}\hat{p}_{t}(x)\Big{[}e_{\ell}(x)\left(1-\hat{\mu}_{t}(x,1)\right)+(1-e_{\ell}(x))\left(1-\hat{\mu}_{t}(x,0)\right)\Big{]}.$
@@
In the following sections, we will provide theoretical guarantees and simulation evidence to demonstrate the advantages of our proposed fully forward-looking CARA designs in the presence of delayed outcomes.
@@ 4 Theoretical investigation
This section provides theoretical justifications for the proposed forward-looking CARA design. To begin, we present the semiparametric efficiency bound for treatment effect estimation in a multi-stage setting with delayed outcomes. We then develop a general result on the consistency and asymptotic normality of the estimated average treatment effect. Our theoretical investigation then proceeds by examining the specific setting considered in Section 3, where we extrapolate the estimated delay mechanism in each stage using a conservative rule. Finally, we observe that in an experiment with only finitely many stages, it is generally not possible to achieve the oracle allocation, as this would require perfect ex ante knowledge of the delay mechanism. Therefore, the last part of our theoretical investigation addresses the efficiency loss resulting from extrapolating the delay mechanism in our design method.
@@ 4.1 Notation, assumptions and the efficiency bound
We note that while the previous theorem is developed in the context of discrete covariates (with the efficiency bound expressed as a summation over the probability mass function), it can be generalized to accommodate, for example, continuously distributed covariates (c.f., equation 1).
@@ 4.2 Asymptotic properties of treatment effects estimates
In this subsection, we demonstrate that the treatment effect estimator is consistent and admits an asymptotic normal distribution. The main conclusions of this section rely on a high-level condition regarding the convergence of the optimized treatment allocation rule, making these conclusions applicable to a wide range of CARA and a broad class of frequentist adaptive experimental design settings (Hu and Zhang, 2004; Antognini and Zagoraiou, 2015; Hu, Zhu and Hu, 2015). We begin with this high-level condition, which will be verified in the next subsection for our proposed forward-looking CARA design.
@@
To conclude this subsection, the following result establishes valid statistical inference, where the variance estimator $\hat{\mathsf{V}}_{T}$ is defined at the end of Section 3.
@@ 4.3 Convergence of the optimized treatment allocation
To start, we recall that $\mathsf{V}(e_{1}(\cdot),\dots,e_{T}(\cdot))$ is the objective function in (3). Its minimizers will be denoted by $e_{t}^{*}(a|x)$ for $a\in\{0,1\}$, $x\in\mathcal{X}$, and $t=1,2,\dots,T$. As we have discussed, the oracle optimal solution is not achievable in general, as it builds on perfect knowledge of the delay mechanism. In any realistic experimental setting, however, this information can only be learned sequentially. In fact, even we set $n_{t}=\infty$, part of the delay mechanism can still remain unknown for any $t$: to be even more precise, one typically only learn $\rho(d|x,a)$ for $d<t$ using information at the end of stage $t$. For this reason, we have introduced feasible and concrete ways to extrapolate the delay mechanism. Following our discussion in Section 2, we define
$\rho^{\dagger}_{t}(d|x,a)$ $=\min\big{\{}\rho(d|x,a),\rho(t-1|x,a)\big{\}}.$
@@ 5 Synthetic case study
To evaluate the performance of the proposed forward-looking CARA design, we conduct simulation studies using a data generation mechanism calibrated from the study of Fahey, Njau, Katabaro, Mfaume, Ulenga, Mwenda et al. (2020). Their original experiment aims to evaluate the effectiveness of cash transfers on retention in care among people living with HIV (PLHIV), using viral load as the primary response variable. The experiment participants were adults aged 18 years or older diagnosed with HIV, who were randomized to receive cash transfers (treatment arm, $A_{i}=1$) or no cash transfer (control arm, $A_{i}=0$). The original study involved two treatment arms with different levels of cash incentives, which are combined into one group in our synthetic case study. The response variable, viral load, was assessed six months after enrollment. Although the original study was not conducted in an adaptive fashion, adaptation can be applied since the participants were enrolled sequentially. However, a challenge is that the primary responses of the participants were collected with delays, with some responses collected longer than six months after the treatment assignment; see Figure 1. Our synthetic case study conducts a simulation experiment in which the underlying data-generating process (delay mechanism, covariate distribution, and potential outcome distributions) is calibrated with estimates from the original study. For the first set of simulation results, we consider biological sex as the only covariate $X_{i}$, leading to two subgroups: $X_{i}=\mathtt{M}$ and $X_{i}=\mathtt{F}$ for male and female participants. From the original data, the distribution of biological sex consists of 36% male and 64% female, which we adopt as the data-generating process of $X_{i}$ in our Monte Carlo exercise. We set $n_{t}=100$ for $t=1,\ldots,T$ with $T\in\{4,6,8\}$. We report simulation evidence separately for the two design objectives in (3) and (4), which require different transformation of the outcome variable.
@@
Setup 1: Power maximization. For the power maximization design objective, we compare four design methods: (i) our proposed design; (ii) our design enhanced by the doubly adaptive biased coin design (DBCD) (Hu and Zhang, 2004; Tymofyeyev, Rosenberger and Hu, 2007); (iii) complete randomization; (iv) Neyman allocation. We note that DBCD is a response-adaptive randomization design that more robustly adjusts treatment allocation toward the optimal allocation dynamically. Since our approach involves assigning treatments across multiple covariate strata, we apply the DBCD design separately within each stratum. We also follow the convention in the literature and use the natural logarithm of the viral load as the outcome variable. Following the trial data, the conditional mean and standard deviation of the potential outcomes are $\mu(\mathtt{F},1)=2.50$, $\mu(\mathtt{F},0)=2.98$, $\sigma(\mathtt{F},1)=0.36$, $\sigma(\mathtt{F},0)=2.06$ for the female subgroup, and $\mu(\mathtt{M},1)=2.47$, $\mu(\mathtt{M},0)=2.72$, $\sigma(\mathtt{M},1)=0.82$, $\sigma(\mathtt{M},0)=0.31$ for the male subgroup. Given the calibrated model primitives, the average treatment effect is $\tau=-0.33$. To evaluate the power of different designs under the power maximization objective, we also vary the magnitude of the ATE in our data-generating process by perturbing $\mu(\mathtt{M},0)$, such that the perturbed ATE takes values between 0 and $0.66$.
@@
Setup 2: Failure reduction. For the failure reduction objective, we again compare across four designs: (i) the proposed method; (ii) our proposed design again enhanced by DBCD; (iii) the ethical design proposed by Rosenberger, Stallard, Ivanova, Harper and Ricks (2001); (iv) complete randomization. In this simulation setup, we transform the original viral load measurements into a binary variable, indicating whether the participant has achieved viral suppression; specifically, if a participant’s viral load is below $1,000$ copies per mL, we set $Y_{i}=1$ (viral suppression). Conditional means of the potential outcomes are $\mu(\mathtt{F},1)=0.78$, $\mu(\mathtt{F},0)=0.57$, $\mu(\mathtt{M},1)=0.84$ and $\mu(\mathtt{M},0)=0.63$. The average treatment effect is $\tau=0.21$. To evaluate the power of different designs under the failure reduction objective, we vary the magnitude of ATE in our data-generating process by perturbing $\mu(\mathtt{M},0)$, such that the perturbed ATE takes values between 0 and $0.42$.
@@
To assess the finite-sample distributional properties of the treatment effect estimator, we present the coverage probabilities in Table 1 for the two design objectives and the three extrapolations methods for the estimated delay mechanism. Encouragingly, the empirical coverage is extremely close to the nominal level. Lastly, we demonstrate and compare the power properties of different design strategies in Figure 4: all methods control the type I error rate well when the true ATE is zero; with $\tau>0$, however, our fully forward-looking CARA design has a clear advantage as it is able to detect the non-zero treatment effect more frequently.
In summary, this synthetic case study showcases that our proposed fully forward-looking CARA design not only provides valid statistical inference in the presence of delayed responses, but it also delivers higher estimation efficiency and enhanced failure reduction for the two design objectives, which is in stark contrast to other methods that overlook the outcome delay issue.
@@
Before closing this section, we provide additional simulation evidence with a different covariate stratification, defined by biological sex and the WHO clinical stages of HIV. Specifically, we define four groups: $\texttt{S}_{1}$ for male subjects with WHO stages 1 and 2, $\texttt{S}_{2}$ for female subjects with stages 1 and 2, $\texttt{S}_{3}$ for male with stages 3 and 4, and $\texttt{S}_{4}$ for female with stages 3 and 4.
@@
We present simulation results on the optimized variances in panels (A)–(C) of Figure 5, and the failure rates in panels (D)–(F). Although the covariate stratification differs (cf. Figure 3), a consistent pattern emerges: both our proposed design and its DBCD-enhanced variant yield lower variance and enhanced failure reduction by explicitly accounting for the delay mechanism. To compare, the variances of the treatment effect estimator employing complete randomization are 43.96, 45.17, and 47.82, and the failure rates are 0.34, 0.32, and 0.34, respectively. To complement our earlier findings in Figure 4, we also report power comparisons in Figure 6.

### v3

@@ 1.1 Motivation and Contribution
In clinical trials, social science field experiments, and A/B tests, participants often enroll sequentially, and responses to treatments can vary due to individual-specific characteristics. In these contexts, covariate-adjusted response-adaptive (CARA) designs have gained increasing popularity: by dynamically adjusting treatment assignments based on accumulated covariates and outcome information during the experiment, CARA designs can be adopted to optimize treatment allocation to improve statistical power (Hahn et al., 2011; Hu et al., 2015; Blackwell et al., 2023), reduce participant exposure to less effective treatments and enhance overall welfare (Hu and Rosenberger, 2006; Hu et al., 2015; Wei et al., 2023, 2025), and maintain statistical validity for analyzing experimental data (Zhang et al., 2007; Hu and Hu, 2012; Bugni et al., 2018; Offer-Westort et al., 2021; Bai et al., 2022; Robertson et al., 2023; Zhao, 2023; Wei et al., 2025; Bibaut and Kallus, 2025).
@@
As an example, we revisit the study by Fahey et al. (2020), a randomized experiment conducted in the Shinyanga region of Tanzania between April 24 and December 14, 2018. The study aimed to assess the impact of cash incentives on retention in care and viral suppression among people living with HIV (PLHIV). A total of 530 PLHIV aged 18 or older were sequentially enrolled and randomly assigned to either a treatment group receiving cash incentives or a control group without any incentive. Viral load was measured six months after treatment assignment through blood draws. However, delays occurred, as the lab test required a separate clinical visit and some participants missed their scheduled appointments. We calculated the number of days delayed as the difference between the date of the viral load test and the six-month follow-up date, separately for two covariate strata defined by biological sex. From Figure 1, it is clear that the distribution of outcome delays varies significantly depending on treatment assignment and gender. Specifically, the treatment arm exhibits less delay, and within the treatment arm, the delayed response issue seems less pronounced among female participants. Overall, the empirical observation highlights the importance of accounting for delayed outcomes that are both arm- and covariate-specific.
@@
To our knowledge, although there has been prior work on developing valid statistical inference procedures in the presence of delayed responses in CARA (Bai et al., 2002; Hu et al., 2008; Robertson et al., 2023), no CARA designs has incorporated the delay mechanism when optimizing experimental objectives. In this paper, we address the challenges that response delays pose to covariate-adjusted response-adaptive (CARA) designs. We begin by examining whether existing optimal allocation strategies remain valid under delayed outcomes and, when they do not, derive the optimal allocation in such settings. We then develop methods to adapt CARA designs accordingly. Finally, we propose an efficient estimator that accounts for delayed responses at the end of the experiment. We elaborate on these contributions in more detail below.
@@
First, we demonstrate that the classical optimal allocation strategy used in existing CARA designs no longer provides optimal treatment assignment when accounting for arm- and covariate-dependent delays. Specifically, we first derive a semiparametric efficiency bound for treatment effect estimation in the presence of outcome delay (Theorem 1), which differ substantially from existing results in the literature (Hahn, 1998; Hirano et al., 2003; Cattaneo, 2010; Zhu and Zhu, 2023; Wu et al., 2025). See van der Vaart (2000) and Tsiatis (2006) for general discussions on semiparametric efficiency. This result allows us to show that the optimal delay mechanism involves a unique trajectory of treatment assignment probabilities that vary across different stages, resulting in treatment allocations with lower variance for power maximization, and enhanced failure reduction in welfare improvement designs. To our knowledge, this is the first design to explicitly explore delay mechanisms to optimize trial objectives, while existing literature focuses on conducting sequential tests (Hampson and Jennison, 2013), designing Bayesian adaptive designs where delay is independent of the arm and covariates (Lin et al., 2020), or conducting statistical inference after experiments have concluded (Bai et al., 2002; Hu et al., 2008).
@@
Second, since the delay mechanism is unknown a priori and remains only partially estimable during the experiment, we propose a fully forward-looking CARA design that sequentially updates treatment assignment probabilities to target a trial objective at the conclusion of the experiment. In this process, our algorithm offers several extrapolation strategies when estimating delay mechanism, allowing better alignment with the designer’s prior knowledge. We further discuss in detail two design objectives frequently adopted in current CARA literature: maximizing experiment power (Tymofyeyev et al., 2007; Zhao, 2023) and improving overall participant welfare (or failure reduction) while maintaining statistical power constraints (Rosenberger et al., 2001; Robertson et al., 2023). We justify the benefits of our proposed forward-looking CARA design through theoretical investigation (Section 4) and a synthetic case study (Section 5). Under the design objective of maximizing experiment power, our proposed forward-looking CARA design attains higher estimation efficiency compared to the Neyman allocation and complete randomization designs in the presence of delayed responses. Under the design objective of improving overall participant welfare, we demonstrate that our proposed forward-looking CARA design achieves additional failure reduction, leading to greater welfare improvement compared to both the ethical design and the complete randomization design while maintaining statistical power. For both design objectives, our proposed design can provide valid statistical inference, which is justified in Theorem 3.
@@ 2 Optimal Delay-adjusted Treatment Allocation in CARA
From the above result, it is evident that when the delay mechanism of the response depends on the covariates and treatment status, the optimal treatment allocation rules in classical CARA designs must be modified to account for outcome delays. In what follows, we examine how the optimal treatment allocation rule is affected by the arm- and/or covariate-dependent delayed responses in two common experimental objectives: one aimed at enhancing statistical power (Tymofyeyev et al., 2007), and the other focused on minimizing the expected number of failures while maintaining a minimum power requirement for binary outcomes (Rosenberger et al., 2001). In both design goals, the variance of the estimated treatment effects plays an essential role. To simplify the presentation, we focus on two-arm CARA designs with $a\in\{0,1\}$ in this paper. For multi-arm experiments, the design objective can be generalized using non-centrality parameters, as discussed in Lachin (1977) and Tymofyeyev et al. (2007).
@@
where $\delta$ ensures that the assignment probabilities are bounded away from 0 and 1, maintaining sufficient randomness in treatment assignments (Ma and Wang, 2020; Heiler and Kazak, 2021; Sasaki and Ura, 2022; Ma et al., 2025a; Dorn, 2025). The objective function $\mathsf{V}(e_{1}(\cdot),\dots,e_{T}(\cdot))$ is a component of the semiparametric efficiency bound related to the propensity scores, and is defined as:
@@
Design objective 2: Optimal treatment allocation for failure reduction subject to a power constraint. This design aims to minimize the expected number of failures while satisfying a minimum power requirement for binary responses (Rosenberger et al., 2001). In this context, a treatment is considered to fail when the potential outcome equals one. The optimal allocation thus seeks to assign treatments based on the solution to the following optimization problem:
@@
and the failure reduction rule of Rosenberger et al. (2001):
@@ 3 Fully Forward-looking CARA with Delayed Outcomes
The results provided in the previous section assume prior knowledge of the delay mechanism, and therefore, the discussed optimal treatment allocation corresponds to an infeasible oracle setting. In practice, the delay mechanism is rarely known and must be estimated as data are sequentially accumulated to optimize the experimental objective. In this section, we propose two delay-informed practical CARA designs tailored to two previously discussed experimental goals: improving statistical power (Tymofyeyev et al., 2007) and minimizing the expected number of failures while meeting a minimum power requirement for binary responses (Rosenberger et al., 2001). We also provide tools for valid statistical inference after the experiment is finished.
@@
$\hat{\mathsf{V}}_{t}(e_{t+1}(\cdot),\dots,e_{T}(\cdot))$ $=\sum_{x\in\mathcal{X}}\hat{p}_{t}(x)\Bigg[\frac{\hat{\sigma}_{t}^{2}(x,1)}{\sum\limits_{\ell=1}^{t}\hat{r}_{\ell}\hat{\rho}_{t}(T-\ell|x,1)\hat{e}^{*}_{\ell}(1|x)+\sum\limits_{\ell=t+1}^{T}\hat{r}_{\ell}\hat{\rho}_{t}(T-\ell|x,1)e_{\ell}(x)}$
$\qquad\qquad+\frac{\hat{\sigma}_{t}^{2}(x,0)}{\sum\limits_{\ell=1}^{t}\hat{r}_{\ell}\hat{\rho}_{t}(T-\ell|x,0)\hat{e}^{*}_{\ell}(0|x)+\sum\limits_{\ell=t+1}^{T}\hat{r}_{\ell}\hat{\rho}_{t}(T-\ell|x,0)(1-e_{\ell}(x))}\Bigg].$
@@
$\quad+\sum_{\ell=t+1}^{T}\hat{r}_{\ell}\sum_{x\in\mathcal{X}}\hat{p}_{t}(x)\Big[e_{\ell}(x)\left(1-\hat{\mu}_{t}(x,1)\right)+(1-e_{\ell}(x))\left(1-\hat{\mu}_{t}(x,0)\right)\Big].$
@@ 4 Theoretical Investigation
This section provides theoretical justifications for the proposed CARA design. To begin, we present the semiparametric efficiency bound for treatment effect estimation in a multi-stage setting with delayed outcomes. We then develop a general result on the consistency and asymptotic normality of the estimated average treatment effect. Our theoretical investigation then proceeds by examining the specific setting considered in Section 3, where we extrapolate the estimated delay mechanism in each stage using a conservative rule. Finally, we observe that in an experiment with only finitely many stages, it is generally not possible to achieve the oracle allocation, as this would require perfect ex ante knowledge of the delay mechanism. Therefore, the last part of our theoretical investigation addresses the efficiency loss resulting from extrapolating the delay mechanism in our design method.
@@ 4.1 Notation, Assumptions and the Efficiency Bound
We note that while the previous theorem is developed in the context of discrete covariates, it can be generalized to accommodate continuously distributed covariates (c.f., equation 1).
@@ 4.2 Asymptotic Properties of Treatment Effects Estimates
In this subsection, we demonstrate that the treatment effect estimator is consistent and admits an asymptotic normal distribution. The main conclusions of this section rely on a high-level condition regarding the convergence of the optimized treatment allocation rule, making these conclusions applicable to a wide range of CARA and a broad class of frequentist adaptive experimental design settings (Hu and Zhang, 2004; Antognini and Zagoraiou, 2015; Hu et al., 2015). We begin with this high-level condition, which will be verified in the next subsection for our proposed forward-looking CARA design.
@@ 4.3 Convergence of the Optimized treatment Allocation
To start, we recall that $\mathsf{V}(e_{1}(\cdot),\dots,e_{T}(\cdot))$ is the objective function in (3). Its minimizers will be denoted by $e_{t}^{*}(a|x)$ for $a\in\{0,1\}$, $x\in\mathcal{X}$, and $t=1,2,\dots,T$. As we have discussed, the oracle optimal solution is not achievable in general, as it builds on perfect knowledge of the delay mechanism. In any realistic experimental setting, however, this information can only be learned sequentially. In fact, even we set $n_{t}=\infty$, part of the delay mechanism can still remain unknown for any $t$: to be even more precise, one typically only learn $\rho(d|x,a)$ for $d<t$ using information at the end of stage $t$. For this reason, we have introduced feasible and concrete ways to extrapolate the delay mechanism. Following our discussion in Section 2, we define $\rho^{\dagger}_{t}(d|x,a)=\min\big\{\rho(d|x,a),\rho(t-1|x,a)\big\}.$
@@ 5 Synthetic Case Study
To evaluate the performance of the proposed forward-looking CARA design, we conduct simulation studies using a data generation mechanism calibrated from the study of Fahey et al. (2020). For the first set of simulation results, we consider biological sex as the only covariate $X_{i}$, leading to two subgroups: $X_{i}=\mathtt{M}$ and $X_{i}=\mathtt{F}$ with frequencies 36% and 64%, respectively. We set $n_{t}=100$ for $t=1,\ldots,T$ with $T\in\{4,6,8\}$. We report simulation evidence separately for the two design objectives in (3) and (4), which require different transformation of the outcome variable.
@@
Setup 1: Power maximization. For the power maximization design objective, we compare four design methods: (i) our proposed design; (ii) our design enhanced by the doubly adaptive biased coin design (DBCD) (Hu and Zhang, 2004; Tymofyeyev et al., 2007); (iii) complete randomization; (iv) Neyman allocation. We note that DBCD is a response-adaptive randomization design that adjusts treatment allocation toward the optimal allocation dynamically. Since our approach involves assigning treatments across multiple covariate strata, we apply the DBCD design separately within each stratum. We also follow the convention in the literature and use the natural logarithm of the viral load as the outcome variable. Following the trial data, the conditional mean and standard deviation of the potential outcomes are $\mu(\mathtt{F},1)=2.50$, $\mu(\mathtt{F},0)=2.98$, $\sigma(\mathtt{F},1)=0.36$, $\sigma(\mathtt{F},0)=2.06$ for the female subgroup, and $\mu(\mathtt{M},1)=2.47$, $\mu(\mathtt{M},0)=2.72$, $\sigma(\mathtt{M},1)=0.82$, $\sigma(\mathtt{M},0)=0.31$ for the male subgroup. Given the calibrated model primitives, the average treatment effect is $\tau=-0.33$. To evaluate the power of different designs under the power maximization objective, we also vary the magnitude of the ATE in our data-generating process by perturbing $\mu(\mathtt{M},0)$, such that the perturbed ATE takes values between 0 and $0.66$.
@@
Setup 2: Failure reduction. For the failure reduction objective, we again compare across four designs: (i–ii) the proposed method (with and without DBCD enhancement); (iii) the ethical design (Rosenberger et al., 2001); (iv) complete randomization. In this simulation setup, we transform the original viral load measurements into a binary variable, indicating whether the participant has achieved viral suppression; specifically, if a participant’s viral load is below $1,000$ copies per mL, we set $Y_{i}=1$ (viral suppression). Conditional means of the potential outcomes are $\mu(\mathtt{F},1)=0.78$, $\mu(\mathtt{F},0)=0.57$, $\mu(\mathtt{M},1)=0.84$ and $\mu(\mathtt{M},0)=0.63$. The average treatment effect is $\tau=0.21$. To evaluate the power of different designs under the failure reduction objective, we vary the magnitude of ATE in our data-generating process by perturbing $\mu(\mathtt{M},0)$, such that the perturbed ATE takes values between 0 and $0.42$.
@@
To assess the finite-sample distributional properties of the treatment effect estimator, we present the coverage probabilities in Table 1 for the two design objectives and the three extrapolations methods for the estimated delay mechanism. Encouragingly, the empirical coverage is extremely close to the nominal level. Lastly, we demonstrate and compare the power properties of different design strategies in Figure 4: all methods control the type I error rate well when the true ATE is zero; with $\tau>0$, however, our CARA design has a clear advantage as it is able to detect the non-zero treatment effect more frequently.
In summary, this synthetic case study showcases that our fully forward-looking CARA design not only provides valid statistical inference in the presence of delayed responses, but it also delivers higher estimation efficiency and enhanced failure reduction for the two design objectives.
@@
Before closing this section, we provide additional simulation evidence where stratification is based on biological sex and the WHO clinical stages of HIV. Specifically, we define four groups: $\texttt{S}_{1}$ for male subjects with WHO stages 1 and 2, $\texttt{S}_{2}$ for female subjects with stages 1 and 2, $\texttt{S}_{3}$ for male with stages 3 and 4, and $\texttt{S}_{4}$ for female with stages 3 and 4.
@@
We present simulation results on the optimized variances in panels (A)–(C) of Figure 5, and the failure rates in (D)–(F). Although the covariate stratification differs (cf. Figure 3), a consistent pattern emerges: our proposed design yields lower variance and enhanced failure reduction by explicitly accounting for the delay mechanism. To compare, the variances of the treatment effect estimator employing complete randomization are 43.96, 45.17, and 47.82, and the failure rates are 0.34, 0.32, and 0.34, respectively. To complement our earlier findings in Figure 4, we also report power comparisons in Figure 6.